# Implementing Adaptations for Vision AutoRegressive Model

### v1

@@ 2 Background and Related Work
where clip$(g(x_{k}))=g(x_{k})/max(1,\frac{||g(x_{k})||_{2}}{C})$, $\eta$ is the learning rate, $C$ is the clipping norm, $\sigma$ is the noise magnitude, $g(x_{k})$ is the gradient for a single input data point $x_{k}$, $L$ is the lot size and $\theta$ are the parameters.
@@ A.2 Differentially Private Diffusion Models
Ghalebikesabi et al. [2023] show that by sampling an augmentation in addition to the timestep at each repetition in the noise multiplicity mechanism yields an improvement. They dub that approach augmentation multiplicity. They find that $K=128$ yields the best results, with batch size of $16,384$. Similarly to DPDM, they sample the timesteps from a distribution different than the default $\mathcal{U}[0,T]$, $t\sim\sum_{i=1}^{K}w_{i}\mathcal{U}[l_{i},u_{i}],\quad\text{where }\sum_{i=1}^{K}w_{i}=1,\,0\leq l_{0},u_{K}\leq T,\,\text{and }u_{k-1}\leq l_{k}\,\text{for}k\in\{2,\dots,K\}$. The main difference between their work and DPDM is that they first pre-train the DM using ”public” dataset without DP, and only then fine-tune on the private dataset, this time with DP.
@@
Liu et al. [2024] focus on the impact of the size of the DM on the properties of DP-SGD. Intuitively, the bigger the size of the model, the bigger the norm of the gradient, in effect the bigger the noise added by the DP-SGD mechanism. Thus, a natural idea is to limit the size of the DM. To this end, authors utilize Latent Diffusion Models (LDMs) [Rombach et al., 2022], which consist of of a VAE and a DM. VAEs are made of an encoder, which takes a high-dimensional image in the pixel space as an input, and produces a lower-dimensional latent representation, and a decoder that converts the latent back to the pixel space. Effectively, the DM is trained in the latent space, and in turn contains less parameters than a pixel-space alternative. DP-LDM first trains the VAE (without DP), and, similarly to [Ghalebikesabi et al., 2023], pre-trains the DM on ”public” dataset. To further benefit from the smaller gradient norms, they fine-tune with DP only the (cross-)attention modules of the DM, which amounts to $10\%$ of the parameters.
@@ Appendix B Extended Experimental Details
Table 8: Hyperparameters for DP-finetuning (DP-$\varepsilon-10$) on Oxford Flowers for VAR.
@@ Appendix C Extended Experimental Results for Non-Private Experiments
Model Adaptation Food-101 CUB-200-2011 Oxford Flowers Stanford Cars Oxford-IIIT Pet Trainable Parameters
@@
Recall values dip sharply for Food-101 (max. $\leq 10\%$) and Oxford-IIIT Pet (all methods $<\!5\%$). Both datasets are (i) fine-grained, demanding the model to cover hundreds of subtly different classes, and (ii) out-of-domain for the ImageNet-trained encoder used by the metric, which clusters visually similar species into overly tight neighbourhoods. Under these conditions the k-NN test rejects many legitimate but stylistically rare samples, driving recall down even for FFT.

### v2

@@ 2 Background and Related Work
[…]
@@ A.2 Differentially Private Diffusion Models
Ghalebikesabi et al. [2023] show that by sampling an augmentation in addition to the timestep at each repetition in the noise multiplicity mechanism yields an improvement. They dub that approach augmentation multiplicity. They find that $K=128$ yields the best results, with batch size of 16,38416,38416 , 384. Similarly to DPDM, they sample the timesteps from a distribution different than the default $\mathcal{U}[0,T]$, $t\sim\sum_{i=1}^{K}w_{i}\mathcal{U}[l_{i},u_{i}],\quad\text{where }\sum_{i=1}^{K}w_{i}=1,\,0\leq l_{0},u_{K}\leq T,\,\text{and }u_{k-1}\leq l_{k}\,\text{for }k\in\{2,\dots,K\}$. The main difference between their work and DPDM is that they first pre-train the DM using ”public” dataset without DP, and only then fine-tune on the private dataset, this time with DP.
@@
[…]
@@ Appendix B Extended Experimental Details
[…]
@@ Appendix C Extended Experimental Results for Non-Private Experiments
Model Adaptation Food-101 CUB-200-2011 Oxford Flowers Stanford Cars Oxford-IIIT Pet
Trainable
Parameters
@@
[…]